A 5.0 kg cart on a table is connected to a string that hangs over pulley suspending a 3.0 kg block over the side of the table. Assuming no friction acts on the system , determine the acceleration of the cart and the tension in the string.


1
Expert's answer
2021-01-05T12:03:16-0500
"(M+m)a=mg\\(5+3)a=3\cdot9.8\\a=3.7\frac{m}{s^2}\\T=Ma=5\cdot 3.7=18\ N"